The product of two consecutive even number is 5928 . what is the greater number?
please help me to solve this question

Found 2 solutions by edjones, checkley77:
Answer by edjones(8007)

Let the numbers be x and x+2.
x(x+2)=5928
x^2+2x-5928=0
(x-76)(x+78)=0
x=76
x+2=78 the greater number
.
Ed

Answer by checkley77(12844)

X(X+2)=5,928
X^2+2X-5,928=0
(X+78)(X-76)=0
X+78=0
X=-78 ANS FOR THE SMALLER NUMBER.
-78+2=-76 ANS. FOR THE LARGER NUMBER.
********************
X-76=0
X=76 ANS FOR THE SMALLER NUMBER.
76+2=78 ANS FOR THE LARGER NUMBER.